Given the perimeter of a rectangle and either the length or width, find the unknown measurement using the appropriate formula.
The perimeter is 114 cm. and the width is 27 cm.

Answer by macston(5194)

The perimeter is 114 cm. and the width is 27 cm
L=length; W=Width; P=perimeter
P=2(L+W) Put in the values
114 cm=2(L+27cm) Divide each side by 2.
57 cm=L+27 cm Subtract 27 cm from each side.
30 cm=L ANSWER: Length=30 cm